Jane averaged her five equally weighted test grades and got 60%. This was wrong because she had entered one grade incorrectly. That incorrect score should have been entered as 80%. If she had used the 80%, her correct average for the five tests would have been 70%. What percent did she incorrectly enter instead of 80%?
Found 2 solutions by stanbon, TimothyLamb:
Answer by stanbon(75887)

Jane averaged her five equally weighted test grades and got 60%. This was wrong because she had entered one grade incorrectly. That incorrect score should have been entered as 80%. If she had used the 80%, her correct average for the five tests would have been 70%. What percent did she incorrectly enter instead of 80%?
--------------
Equations:
4correct scores and 1 error score::
(4c + e)/5 = 60
4c + e = 300
-----------------------
5 correct scores::
(4c + 80)/5 = 70
4c + 80 = 350
-------
Solve this system:
4c + e = 300
4c + 80 = 350
-----
80-e = 50
e = 30
-----
Answer:: 30%
-----------
Cheers,
Stan H.

Answer by TimothyLamb(4379)

x = incorrect test score
y = sum of other four correct test scores
---
x + y = 60*5
80 + y = 70*5
---
put the system of linear equations into standard form
---
x + y = 300
y = 270
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = incorrect test score = 30
y = sum of other four correct test scores = 270
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---